factor the polynomials using whatever strategy seems appropriate. State what methods you will use and then demonstrate the methods on your problems, explaining the process as you go. Discuss any particular challenges those particular polynomials posed for the factoring.
w^2+30+81
ax+xc+aw^2+xw^2
10a^2-27ab+5b^2

Answer by stanbon(75887)

Discuss any particular challenges those particular polynomials posed for the factoring.
w^2+30w+81
Use reverse-FOIL; also called the AC Method
AC = 3*27
B = 3+27
-----
Rewrite:
w^2 + 3w + 27w + 81
-----
w(w+3) + 27(w+3)
-----
= (w+3)(w+27)
=========================
ax+xc+aw^2+xw^2
Note: Probably has a typo.
=========================
10a^2-27ab+5b^2
AC =-25*-2
B = -25+-2
Rewrite:
10a^2 -25ab-2ab + 5b^2
-----
5a(2a-5b) - b(2a-5b)
-----
(2a-5b)(5a-b)
=====================
Cheers,
Stan H.
=====================